# Entanglement scaling behaviors of free fermions on hyperbolic lattices

### v1

@@ Abstract
Recently, tight-binding models on hyperbolic lattices (discretized AdS space), have gained significant attention, leading to hyperbolic band theory and non-Abelian Bloch states. In this paper, we investigate these quantum systems from the perspective of quantum information, focusing particularly on the scaling of entanglement entropy (EE) that has been regarded as a powerful quantum-information probe into exotic phases of matter. It is known that on $d$-dimensional translation-invariant Euclidean lattice, the EE of band insulators scales as an area law ($\sim L^{d-1}$; $L$ is the linear size of the boundary between two subsystems). Meanwhile, the EE of metals (with finite Density-of-State, i.e., DOS) scales as the renowned Gioev-Klich-Widom scaling law ($\sim L^{d-1}\log L$). The appearance of logarithmic divergence, as well as the analytic form of the coefficient $c$ is mathematically controlled by the Widom conjecture of asymptotic behavior of Toeplitz matrices and can be physically understood via the Swingle’s argument. However, the hyperbolic lattice, which generalizes translational symmetry, results in inapplicability of the Widom conjecture and thus presents significant analytic difficulties. Here we make an initial attempt through numerical simulation. Remarkably, we find that both cases adhere to the area law, indicating that the logarithmic divergence arising from finite DOS is suppressed by the background hyperbolic geometry. To achieve the results, we first apply the vertex inflation method to generate hyperbolic lattice on the Poincaré disk, and then apply the Haydock recursion method to compute DOS. Finally, we study the scaling of EE for different bipartitions via exact diagonalization and perform finite-size scaling. We also investigate how the coefficient of the area law is correlated to bulk gap and DOS. Future directions are discussed.
@@ I Introduction
It is worth noting that these scaling behaviors are established on the translation-invariant lattices with Euclidean geometry, where the Widom conjecture of Toeplitz matrices is applicable. Then, it is natural to ask what scaling behavior EE will have in systems with the geometries different from the Euclidean geometry. In fact, non-Euclidean geometry is prevalent in natural and artificial systems, playing important roles in both mathematics and physics [15]. Anti-de Sitter (AdS) space, characterized by negative spatial curvature, is widely studied in various fields of physics such as AdS/CFT correspondence, holography of entanglement and tensor-network theory [16, 17, 18]. The hyperbolic lattice, which can be viewed as a discretization of AdS space [19, 20, 21], is of interest in high energy physics [22, 23, 24, 19, 20, 21, 25, 26]. Recently, hyperbolic lattice has been experimentally realized in the photonic and circuit systems [27, 28, 29, 30, 31] and draws more and more attentions in condensed matter physics, such as quantum phase transition, semimetals and topological features induced by the hyperbolic geometry [32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Hyperbolic lattice is highly different from its Euclidean counterpart due to its symmetry and non-Abelian translation group [42, 43, 44]. Remarkably, these interesting properties lead to higher-dimensional Brillouin zone and hyperbolic band theory (HBT) for tight-binding models on hyperbolic lattices [42, 43, 44].
@@
Motivated by the rapid progress on hyperbolic lattices as well as application of quantum information in many-body physics, we explore the potential role of hyperbolic geometry in affecting quantum entanglement in this paper. However, compared to Euclidean translation-invariant lattice, the analytic difficulties here are significantly challenging as the Widom conjecture is no longer applicable here. Therefore, our goal is to provide numerical evidence of the exotic interplay of quantum entanglement and hyperbolic geometry by investigating the scaling of EE of free-fermion systems on hyperbolic lattices. We observe that for gapped systems, the EE still scales as the area law, consistent with our expectations on Euclidean lattice. However, for gapless system with finite DOS, we discover that the super-area law breaks down, and the EE adheres to the area law instead. The numerical evidence of the violation of super-area law reflects the exotic behavior of free fermions on hyperbolic lattices induced by their geometry. It is worth noting that recent research shows that in free-fermion systems on fractal lattices, the EE of gapless states with finite Density-of-States (DOS) still satisfies the Gioev-Klich-Widom scaling, while the EE of gapped states becomes a generalized area law [45].
@@
This paper is arranged as follows: In Sec. II, we specify the construction of hyperbolic lattices and provide a brief summary of studying free-fermion entanglement entropy. Next in Sec. III we study EE of gapless free-fermion systems with finite DOS and the dependence of scaling coefficient on DOS while in Sec. IV, we study EE of gapped free fermions on hyperbolic lattices. Finally, we summarize our findings in Sec. V and discuss their potential applications. Additionally, we detail the hyperbolic lattice setup in Appendix A, provide supplemental data of EE in Appendix B and review the approach to compute DOS in Appendix C. We also discuss the asymptotic behavior of the coefficient of the area law in Appendix D.
@@ II.4 Entanglement entropy and Widom conjecture
where $\partial\Gamma$ and $\partial\Omega$ denote the boundaries of the Fermi surface and the subsystem we consider, $\bm{n_{p}}$ and $\bm{n_{r}}$ denote the exterior unit normals of these boundaries. Since the presence of codimension-1 Fermi surface implies finite DOS of the system, Eq. (5) also relates the DOS to the scaling coefficient. If the codimension of the Fermi surface is higher than one, the leading term of EE exhibits area law scaling behavior, as seen in the Dirac point of tight-binding model on the honeycomb lattice [7, 66, 67]. However, the validity of Eq. (5) requires a Euclidean metric with Abelian translation symmetry and thus is not naturally applicable in the hyperbolic geometry, so we aim to provide numerical evidence in this paper.
@@ III.1 Numerical study of DOS
Figure 5: Dependence of coefficient $a$ on DOS of $\{4,5,6\}$ lattice, computed for Hamiltonian $H_{1}$ with $t=1$ and $\mu\in[0,4]$. The $p$-value approaching 00 and $\rho$ approaching $1$ of the Spearman’s correlation verifies the positive relation between scaling coefficient and DOS, suggesting that a generalized Widom conjecture may exist.
@@ III.2 Numerical evidence of area law scaling behavior of EE
To proceed further, we use our approaches detailed in Sec. II to compute EE on various lattices, including both Euclidean and hyperbolic. In Fig. 4(c2-c4) we show the results computed on $\{3,7\}$, $\{4,5\}$, $\{6,4\}$ lattices. Additionally, in Fig. 4(c1), we also include EE computed on Euclidean $\{4,4\}$ lattice for comparison. More numerical results through different partition methods on different lattices are detailed in Appendix B.
First, in the Euclidean case, the EE of gapless systems with finite DOS exhibits super-area law, corresponding to our results computed on $\{4,4\}$ lattice in Fig. 4(c1), where we anticipate the scaling function $S_{A}/\log L_{A}=cL_{A}^{\alpha}+d$. When we turn to the hyperbolic case, our most surprising finding is that the EE of gapless systems with finite DOS is proportional to the length of the boundary of subsystem $A$. We anticipate that the scaling of EE should have $S_{A}=aL_{A}^{\alpha}+b$. By using the coefficient of determination $R^{2}$, we find that $\alpha\approx 1$ is the optimal fit closest to $1$, as shown in Fig. 4(c2-c4). The blue lines in Fig. 4(c) show the fitting functions with $\alpha=1$. This result indicates that the EE of gapless systems with finite DOS on hyperbolic lattices unexpectedly satisfies the area law:
@@
where $L_{A}$ represents the total number of bonds connecting a site inside the subsystem to a site outside the subsystem which are cut by the boundary of the subsystem $A$, as visualized in Fig. 3.
@@
With more numerical computations, as illustrated in Appendix B, we further confirm the existence of area law of EE for gapless ground states with finite DOS on hyperbolic lattice. Furthermore, we want to ask why this exotic area law of EE appears in hyperbolic systems. On Euclidean lattice, following Swingle’s (mode-counting) argument [14], for free-fermion systems with codimension-1 Fermi surface, EE can be obtained by counting the contributions of $1$d fermionic gapless modes near the Fermi surface perpendicular to the boundary of the subsystem in real space, where each fermionic gapless mode contributes $\log L_{A}$ to EE by adopting the calculation of CFT. Then, we obtain that EE satisfies $S_{A}\sim L_{A}^{d-1}\log L_{A}$ in Euclidean case. On hyperbolic lattices, the gapless system with finite DOS still has infinite gapless fermionic modes near Fermi level despite the absence of “Fermi surface” of the usual definition. If we can stack and count the contribution of these infinite fermionic gapless modes for EE, we can obtain the scaling behavior of EE for hyperbolic systems. However, there is a lack of a realizable stacking and counting way on hyperbolic lattice. According to Eq. (4), EE depends on the projectors $\hat{P}$ and $\hat{R}$. HBT provides an insight for us into the parameterization of the generalized hyperbolic Brillouin zone and non-Abelian Bloch states [42, 43, 35, 44, 33]. Therefore, our numerical simulation raises questions and challenges for HBT to obtain a generalized Widom conjecture for hyperbolic lattices, as well as the expressions of $\hat{P}$ and $\hat{R}$ from the parameterized momentum space.
@@
Moreover, the scaling behavior of EE is related to the non-local properties of the systems; therefore, the super-area law behavior of EE indicates that fermionic statistics enhances entanglement in Euclidean geometry. Due to the absence of the logarithmic correction of EE in Eq. (9), we realize that the gapless fermions in hyperbolic lattices should have exotic behavior. Additionally, as hyperbolic geometry suppress entanglement, it is worth investigating the asymptotic behavior of EE with respect to $q$ and we discuss this in Appendix D. In the forthcoming Sec. III.3, we will continue to discuss our numerical findings, especially focusing on the scaling coefficient.
@@ III.3 Numerical study of scaling coefficient and possibility of a generalized Widom conjecture
In the Euclidean case, we know from Eq. (5) that the scaling coefficient of the super-area law is determined by the geometry of the codimension-1 Fermi surface in momentum space and the partition of subsystems in real space, and thus is related to DOS, as visualized in Fig. 4(d1). From this perspective, we question whether DOS can affect the coefficient $a$ in area law $S_{A}=aL_{A}+\cdots$ for systems on hyperbolic lattice, and thus consider its dependence on the DOS.
We compute EE for Hamiltonian $H_{1}$ with $t=1$ and different chemical potential $\mu$ on different hyperbolic lattices. Fig. 4(d2-d4) shows the dependence of the scaling coefficient $a$ on $\mu$. Compared to the DOS computed in Fig. 4(b2-b4), we can directly see that the scaling coefficient $a$ is positively related to DOS. In Fig. 5, we present the result computed on $\{4,5,6\}$ lattice as an example, where the Spearman’s correlation indicates a positive monotonic relationship between $a$ and the DOS. However, from Fig. 4, we can see that $a$ and the DOS do not completely coincide. This discrepancy might also be due to the finite-size effect, as the EE computed here is obtained from a finite lattice while the DOS is obtained in the thermodynamical limit.
In the Euclidean system, the validity of the Widom conjecture needs a Euclidean metric and the momentum space with dimension equal to real-space dimension due to the flux factor $\left|\bm{n_{r}}\cdot\bm{n_{p}}\right|$ in Eq. (5) counting the number of fermionic modes perpendicular to the real space boundary of the subsystem. For the hyperbolic case, since scaling coefficient $a$ of area law exhibits similar dependence on DOS, we can also speculate whether there is a similar momentum space and generalized Fermi surface that determines $a$. In fact, the translation group of hyperbolic lattice is typically non-Abelian, resulting in the existence of higher-dimensional ($d\geq 2$) irreducible representations of translation group and non-Abelian Bloch states. Meanwhile, even for the $1$d irreducible representations, the dimension of the generalized Brillouin zone can be $d>2$ [42, 43, 44], which is larger than the spatial dimension of the lattice, thus Swingle’s argument breaks down directly. To exactly obtain a description of reciprocal space of hyperbolic lattice, one needs to know about the higher-dimensional representations. It is an open question that whether we can obtain a generalized Widom conjecture for hyperbolic lattice.

### v2

@@ Abstract
Recently, tight-binding models on hyperbolic lattices (discretized AdS space) have gained significant attention, leading to hyperbolic band theory and non-Abelian Bloch states. In this paper, we investigate these quantum systems from the perspective of quantum information, focusing particularly on the scaling of entanglement entropy (EE) that has been regarded as a powerful quantum-information probe into exotic phases of matter. It is known that on $d$-dimensional translation-invariant Euclidean lattice, the EE of band insulators scales as an area law ($\sim L^{d-1}$; $L$ is the linear size of the boundary between two subsystems). Meanwhile, the EE of metals (with finite Density-of-State, i.e., DOS) scales as the renowned Gioev-Klich-Widom scaling law ($\sim L^{d-1}\log L$). The appearance of logarithmic divergence, as well as the analytic form of the coefficient $c$ is mathematically controlled by the Widom conjecture of asymptotic behavior of Toeplitz matrices and can be physically understood via the Swingle’s argument. However, the hyperbolic lattice, which generalizes translational symmetry, results in inapplicability of these analytic approaches and the potential non-trivial behavior of EE. Here we make an initial attempt through numerical simulation. Remarkably, we find that both cases adhere to the area law, indicating the effect of background hyperbolic geometry that influences quantum entanglement. To achieve the results, we first apply the vertex inflation method to generate hyperbolic lattice on the Poincaré disk, and then apply the Haydock recursion method to compute DOS. Finally, we study the scaling of EE for different bipartitions via exact diagonalization and perform finite-size scaling. We also investigate how the coefficient of the area law is correlated to bulk gap in gapped case and to the DOS in gapless case respectively. Future directions are discussed.
@@ I Introduction
It is worth noting that these scaling behaviors are established on the translation-invariant lattices with Euclidean geometry, where the Widom conjecture of asymptotic behavior of Toeplitz matrices is applicable. Therefore, we raise the question whether the behaviors of EE could be significantly changed by the background geometry, as we find that entanglement on fractal lattice can exhibit fractal-like distribution and generalized area law reflecting boundary Hausdorff dimension [15]. In fact, non-Euclidean geometry is prevalent in natural and artificial systems [16]. Anti-de Sitter (AdS) space, characterized by negative spatial curvature, is widely studied in various fields of physics [17, 18, 19, 20, 21, 22, 23, 24, 25]. The hyperbolic lattice, which can be viewed as a discretization of AdS space, is of interest in high energy physics [26, 27, 28, 29]. Recently, hyperbolic lattice has been experimentally simulated on many platforms [30, 31, 32, 33, 34, 35] and draws more and more attentions in various fields of condensed matter physics [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Hyperbolic lattice is highly different from its Euclidean counterpart due to its non-Abelian translation symmetry [54, 55, 56, 57]. Remarkably, these geometric properties lead to the hyperbolic band theory (HBT) for tight-binding models on hyperbolic lattices [55, 56, 57, 58, 59].
@@
The absence of Euclidean translation invariance on hyperbolic lattice results in the inapplicability of Widom conjecture of Toeplitz matrices, implying that the EE may exhibit non-trivial behaviors. Motivated by the rapid progress on hyperbolic lattices as well as application of quantum information in many-body physics, we explore the potential role of hyperbolic geometry in affecting quantum entanglement in this paper. However, the analytic difficulties are significantly challenging as the Widom conjecture of Toeplitz matrices is no longer applicable. Therefore, our goal is to provide numerical evidence of the exotic interplay of quantum entanglement and hyperbolic geometry by investigating the scaling of EE of free-fermion systems on hyperbolic lattices. We observe that for gapped systems, the EE still scales as the area law, consistent with our expectations on Euclidean lattice. However, for gapless system with finite DOS, we discover that the super-area law breaks down, and the EE adheres to the area law instead. This area law scaling reflects non-trivial effect of hyperbolic geometry for entanglement, which may relates to a holographic understanding that could be experimentally studied [19, 34]. Moreover, similar to our previous work in fractal geometry as summarized in Table. 1, our results suggest a perspective to study the geometry of quantum systems through entanglement.
Table 1: Scaling behavior of EE of ground states of free fermionic systems on $d$-dimensional Euclidean lattice with translation invariance, fractal lattice with self-similarity and two-dimensional hyperbolic lattice. In fractal case, $d_{\rm bf}$ denotes the Hausdorff dimension of boundary of subsystem $A$, while EE of gapless systems exhibits fractal-like distribution [15].
Lattice Phase $S_{A}$
Euclidean lattice [2, 3] Gapped,  DOS$=0$ $\sim L_{A}^{d-1}$
Gapless,  DOS$>0$ $\sim L_{A}^{d-1}\log L_{A}$
Fractal lattice [15] Gapped,  DOS$=0$ $\sim L_{A}^{d_{\rm bf}}$
Gapless,  DOS$>0$ $\sim L^{d-1}_{A}\log L_{A}$
hyperbolic lattice Gapped,  DOS$=0$ $\sim L_{A}$
Gapless,  DOS$>0$ $\sim L_{A}$
@@
This paper is arranged as follows: In Sec. II, we specify the construction of hyperbolic lattices and provide a brief summary of studying free-fermion entanglement entropy. Next in Sec. III we study EE of gapless free-fermion systems with finite DOS and the dependence of scaling coefficient on DOS while in Sec. IV, we study EE of gapped free fermions on hyperbolic lattices. Finally, we summarize our findings in Sec. V and discuss their potential applications. Additionally, we detail the hyperbolic lattice setup and discuss the volume law in Appendix A, provide supplemental data of EE in Appendix B and review the approach to compute DOS in Appendix C. We provide the finite-size scaling analysis in Appendix D and the analysis of super-area law in Appendix E. We discuss the asymptotic behavior of the coefficient of the area law in Appendix F.
@@ II.4 Entanglement entropy and Widom conjecture
where $\partial\Gamma$ and $\partial\Omega$ denote the boundaries of the Fermi sea and the subsystem we consider, $\bm{n_{p}}$ and $\bm{n_{r}}$ denote the exterior unit normals of these boundaries. Since the presence of codimension-1 Fermi surface implies finite DOS of the system, Eq. (5) also relates the DOS to the scaling coefficient. If the codimension of the Fermi surface is higher than one, the leading term of EE exhibits area law scaling behavior, as seen in the Dirac point of tight-binding model on the honeycomb lattice [7, 79, 80]. However, the validity of Eq. (5) requires a Euclidean metric with Abelian translation symmetry and thus is not naturally applicable in the hyperbolic geometry, so we aim to provide numerical evidence in this paper.
@@ III.2 Numerical evidence of area law scaling behavior of EE
To proceed further, we use our approaches detailed in Sec. II to compute EE on various lattices, including both Euclidean and hyperbolic. In Fig. 4(c2-c4) we show the results computed on $\{3,7\}$, $\{4,5\}$, $\{6,4\}$ lattices. Additionally, in Fig. 4(c1), we also include EE computed on Euclidean $\{4,4\}$ lattice for comparison. More numerical results through different partition methods on different lattices are detailed in Appendix B. The finite-size scaling analysis can be found in Appendix D. The numerical analysis of super-area law is presented in Appendix E.
First, in the Euclidean case, the EE of gapless systems with finite DOS exhibits super-area law, corresponding to our results computed on $\{4,4\}$ lattice in Fig. 4(c1), where we anticipate the scaling function $S_{A}/\log L_{A}=cL_{A}^{\alpha}+d$. When we turn to the hyperbolic case, our most surprising finding is that the EE of gapless systems with finite DOS is proportional to the length of the boundary of subsystem $A$. We anticipate that the scaling of EE should have $S_{A}=aL_{A}^{\alpha}+b$. By using the coefficient of determination $R^{2}$, we find that $\alpha\approx 1$ is the optimal fit closest to $1$, as shown in Fig. 4(c2-c4). The blue lines in Fig. 4(c) show the fitting functions with $\alpha=1$. This result indicates that the EE of gapless systems with finite DOS on hyperbolic lattices satisfies the area law by definition:
@@
where $L_{A}$ represents the total number of bonds connecting a site inside the subsystem to a site outside the subsystem which are cut by the boundary of the subsystem $A$, which is consistent with the Euclidean case and visualized in Fig. 3. As the number of sites on the boundary can grow linearly with the number of sites in the subsystem in the thermodynamical limit, we discuss the volume law of EE in Appendix A.
With more numerical computations, as illustrated in Appendix B, and through the numerical analysis that excludes of the possibility of super-area law scaling presented in Appendix E, we further confirm the existence of area law of EE for gapless ground states with finite DOS on hyperbolic lattice. Recent Reference [34] experimentally simulates weakly-coupled scalar field to study AdS/CFT correspondence on hyperbolic lattice. In this work, the EE behavior for entanglement-wedge subsystems of the bulk scalar field satisfies the Ryu-Takayanagi (RT) formula [19] for the connection between the EE of boundary CFT and geometry of the hyperbolic lattice, a result that also has physical understanding [19, 17, 18]. Furthermore, we want to ask why this exotic area law Eq. (9) of gapless free-fermion systems with finite DOS appears in hyperbolic case. The analytical formula of the EE is based on the Widom conjecture of the asymptotic behaviors of Toeplitz matrices on the Euclidean lattice. Due to the absence of Euclidean translation invariance on hyperbolic lattice, the future analytical understanding of EE could be associated with studying conjecture of correlation matrices with symmetry of hyperbolic lattice.
@@
Moreover, following Swingle’s mode-counting argument [14], for free-fermion systems with codimension-1 Fermi surface, EE can be obtained by counting the contributions of $1$d fermionic gapless modes near the Fermi surface perpendicular to the boundary of the subsystem in real space, where each fermionic gapless mode contributes $\log L_{A}$ to EE by adopting the calculation of CFT. Then, we obtain that EE satisfies $S_{A}\sim L_{A}^{d-1}\log L_{A}$ in Euclidean case. On hyperbolic lattice, the Swingle’s mode-counting picture is invalid due to the absence of “Fermi surface” of the usual definition. If we can stack and count the contribution of the infinite fermionic gapless modes near the generalized “Fermi surface” for EE, we can obtain the scaling behavior of EE for hyperbolic systems. However, there is a lack of a realizable stacking and counting way on hyperbolic lattice. According to Eq. (4), EE depends on the projectors $\hat{P}$ and $\hat{R}$. HBT provides an insight for us into the parameterization of the generalized hyperbolic momentum space and non-Abelian Bloch states [55, 56, 57]. Therefore, our numerical simulation raises questions and challenges for HBT to obtain a generalized Widom conjecture and Swingle’s mode-counting picture for hyperbolic lattices, as well as the expressions of $\hat{P}$ and $\hat{R}$ from the parameterized momentum space.
@@
The scaling behavior of EE is related to the non-local properties of the systems. Due to the absence of the logarithmic correction of EE in Eq. (9), we realize that the gapless fermions on hyperbolic lattices should have exotic behavior owing to its non-trivial underlying hyperbolic geometry, and the study of this area law may provide perspective of entanglement for HBT as discussed above. Additionally, as hyperbolic geometry suppresses entanglement, it is worth investigating the asymptotic behavior of EE with respect to $q$ and we discuss this in Appendix F. In the forthcoming Sec. III.3, we will continue to discuss our numerical findings, especially focusing on the scaling coefficient.
Figure 5: Dependence of coefficient $a$ on DOS of $\{4,5,6\}$ lattice, computed for Hamiltonian $H_{1}$ with $t=1$ and $\mu\in[0,4]$. The $p$-value approaching 00 and $\rho$ approaching $1$ of the Spearman’s correlation verifies the approximately positive correlation between scaling coefficient and DOS, suggesting that a generalized Widom conjecture may exist.
In the Euclidean case, we know from Eq. (5) that the scaling coefficient of the super-area law is analytically determined by the flux factor $\left|\bm{n_{r}}\cdot\bm{n_{p}}\right|$ reflecting geometry of the codimension-1 Fermi surface and thus the scaling coefficient changes according to DOS, as visualized in Fig. 4(d1). Because hyperbolic lattice allows for generalized momentum space, we question whether the DOS can influence the coefficient $a$ in area law Eq. (9) following the Euclidean scenario.
We compute EE for Hamiltonian $H_{1}$ with $t=1$ and different chemical potential $\mu$ on different hyperbolic lattices. Fig. 4(d2-d4) shows the dependence of the scaling coefficient $a$ on $\mu$. Compared to the DOS computed in Fig. 4(b2-b4), we can directly see that the scaling coefficient $a$ is correlated to the DOS. In Fig. 5, we present the result of $a$ computed on $\{4,5,6\}$ lattice as an example, where the Spearman’s correlation indicates the approximately positive correlation between $a$ and the DOS. Notably, from Fig. 4, we can see that $a$ and the DOS do not completely coincide. This discrepancy might also be due to the finite-size effect, as the EE computed here is obtained from a finite lattice while the DOS is obtained in the thermodynamical limit.
Remarkably, the scaling coefficient $a$ is non-universal and influenced by many factors such as partition and lattice configuration, but the approximately positive correlation between $a$ and the DOS implies that a generalized momentum space and Fermi surface might play a role in determining $a$, similar to its Euclidean counterpart. In the Euclidean system, the validity of the Widom conjecture and Swingle’s mode-counting picture need a Euclidean metric and the momentum space with dimension equal to real-space dimension due to the flux factor $\left|\bm{n_{r}}\cdot\bm{n_{p}}\right|$ in Eq. (5) counting the number of fermionic modes perpendicular to the real space boundary of the subsystem. In fact, the translation group of hyperbolic lattice is typically non-Abelian, resulting in the existence of higher-dimensional ($d\geq 2$) irreducible representations of translation group and non-Abelian Bloch states. Meanwhile, even for the $1$d irreducible representations, the dimension of the generalized momentum space can be $d>2$ [55, 56, 57], which is larger than the spatial dimension of the lattice, thus Swingle’s argument breaks down directly. To exactly obtain a description of reciprocal space of hyperbolic lattice, one needs to know about the higher-dimensional representations. It is an open question that whether we can obtain a generalized Widom conjecture and Swingle’s mode-counting picture for hyperbolic lattice.